# Simple and Nearly-Optimal Sampling for Rank-1 Tensor Completion via Gauss-Jordan

### v1

@@ 1 Introduction
By contrast, in this work we describe a linear algebraic characterization, which was independently observed by [SSZ20] for cubic tensors. However, the work did not leverage this to study the statistical hardness of our problem. We show Rank-1 Completion admits an extremely simple algorithm, whose main bottleneck is solving a $O(d^{2})\times O(d)$ system over $\mathbb{F}_{2}$. As a consequence, the toolbox for our analysis only consists of elementary linear algebraic and probabilistic arguments.
@@
If $\mathcal{U}\in\otimes_{i=1}^{N}\mathbb{R}_{\neq 0}^{d}$ is a rank-1 tensor with $\|\mathcal{U}\|_{\infty}\leq\rho$ for some $\rho>0$, then there exists an algorithm solving Rank-1 Completion using $m\lesssim(dN)^{2}\log d$ uniformly drawn entries, and can be implemented in time $\lesssim qN+md^{2}$, where $q$ is the number of queried entries of $\mathcal{U}$.
@@ 2 Nonzero Rank-1 Tensors in Exponential Form
$\mathbf{x}:=\begin{pmatrix}\mathbf{u}_{1}\\ \mathbf{u}_{2}\\ \vdots\\ \mathbf{u}_{N}\end{pmatrix}=\begin{pmatrix}\begin{pmatrix}(\mathbf{u}_{1})_{1}\\ \vdots\\ (\mathbf{u}_{1})_{d}\end{pmatrix}\\ \vdots\\ \begin{pmatrix}(\mathbf{u}_{N})_{1}\\ \vdots\\ (\mathbf{u}_{N})_{d}\end{pmatrix}\end{pmatrix}$ (3)
@@
Notably, $\mathbf{A}$’s rows enumerate all $d^{N}$ possible vectors of size $dN$ obtained by concatenating $N$ row vectors from $\{\mathbf{e}_{k}\}_{k\in[d]}$. This observation enables a simple proof of a result we use extensively—that $\mathbf{A}$ has the same rank considered as matrix over $\mathbb{F}_{2}$ or $\mathbb{R}$.
@@
$\mathcal{U}=\operatorname{vec}_{\pi}^{-1}(\varphi^{-1}(\mathbf{A}\mathbf{y}_{1})\odot\exp(\mathbf{A}\mathbf{y}_{2}))$
@@
where the rows adheres to the ordering induced by $\pi$. We similarly define $\mathbf{D}_{\bar{S}}$ for the complement $\bar{S}:=[d]^{N}\setminus S$. As a result, there is always a row permutation matrix $\mathbf{P}_{S}\in\{0,1\}^{d^{N}\times d^{N}}$ where
@@
As we’ve established, since each entry corresponds with a particular row of augmented systems, we have the next observation.
@@
Hence, we can use Algorithm 1 to solve Rank-1 Completion by simply running it for a randomly drawn $S$. Clearly then, we only need to show these conditions holds with sufficient probability. We prove that $O((dN)^{2}\log d)$ samples suffice, although we conjecture it can be improved to $O(dN\log dN)$. Since the result below is crucial for Theorem 1.5, we overview our proof technique.

### v2

@@ 1 Introduction
By contrast, in this work we describe a linear algebraic characterization, which was independently observed by [SSZ20] for cubic tensors. However, the work did not leverage this to study the statistical hardness of our problem. We show Rank-1 Completion admits an extremely simple algorithm, whose main bottleneck is solving a $\tilde{O}(d^{2})\times\tilde{O}(d)$ system over $\mathbb{F}_{2}$. As a consequence, the toolbox for our analysis only consists of elementary linear algebraic and probabilistic arguments.
@@
If $\mathcal{U}\in\otimes_{i=1}^{N}\mathbb{R}_{\neq 0}^{d}$ is a rank-1 tensor with $\|\mathcal{U}\|_{\infty}\leq\rho$ for some $\rho>0$, then there exists an algorithm solving Rank-1 Completion using $m\lesssim(dN)^{2}\log d$ uniformly drawn entries, and can be implemented in time $\lesssim qN+m(dN)^{2}$, where $q$ is the number of queried entries of $\mathcal{U}$.
@@ 2 Nonzero Rank-1 Tensors in Exponential Form
$\mathbf{x}:=\begin{pmatrix}\mathbf{u}_{1}\\ \mathbf{u}_{2}\\ \vdots\\ \mathbf{u}_{N}\end{pmatrix}=\begin{pmatrix}\begin{pmatrix}(\mathbf{u}_{1})_{1}\\ \vdots\\ (\mathbf{u}_{1})_{d}\end{pmatrix}\\ \vdots\\ \begin{pmatrix}(\mathbf{u}_{N})_{1}\\ \vdots\\ (\mathbf{u}_{N})_{d}\end{pmatrix}\end{pmatrix}.$ (3)
@@
Notably, $\mathbf{A}$’s rows enumerate all $d^{N}$ possible vectors of size $dN$ obtained by concatenating $N$ row vectors from $\{\mathbf{e}_{k}\}_{k\in[d]}$. This observation enables a simple proof of a result we use extensively—that $\mathbf{A}$ has the same rank considered as a matrix over $\mathbb{F}_{2}$ or $\mathbb{R}$.
@@
$\mathcal{U}=\operatorname{vec}_{\pi}^{-1}(\varphi^{-1}(\mathbf{A}\mathbf{y}_{1})\odot\exp(\mathbf{A}\mathbf{y}_{2})),$
@@
where the rows adhere to the ordering induced by $\pi$. We similarly define $\mathbf{D}_{\bar{S}}$ for the complement $\bar{S}:=[d]^{N}\setminus S$. As a result, there is always a row permutation matrix $\mathbf{P}_{S}\in\{0,1\}^{d^{N}\times d^{N}}$ where
@@
As we’ve established, since each entry corresponds with a particular row of the augmented systems, we have the next observation.
@@
Hence, we can use Algorithm 1 to solve Rank-1 Completion by simply running it for a randomly drawn $S$. Clearly, we only need to show these conditions hold with sufficient probability. We prove that $O((dN)^{2}\log d)$ samples suffice, although we conjecture it can be improved to $O(dN\log dN)$. Since the result below is crucial for Theorem 1.5, we overview our proof technique.